h(x)e^-x Description of transformation, Equations for the Horizontal Asymptote, y-intercept in (x,y) form. I am not sure how to do this one becuase of the -x? Help
Answer by Nate(3500)

e^-x = 1 / e^x
Transformation: (x,y) ~> (-x,y) so flip about y-axis
Horizontal asymptote is the value for y when x approaches infinity, or increases indefinately.
1 / e = 0.37
1 / e^2 = 0.14
1 / e^3 = 0.05
1 / e^4 = 0.02
Horizontal Asymptote: y = 0
Y-intercept as coordinates (0,y)
y = 1 / e^x
y = 1 / e^0
y = 1
So: (0,1)